Prof. Julio Rezende, PhD
Prof. Julio Rezende, PhD
COMMANDER REPORT: November 7th – MISSION MDRS 182 – 4th Day of Simulation
Yesterday, after conclude the CapComm, I prepared the dinner: eggs with broccoli, tomato soup and tortilla chips. The people liked the tomato soup. J
Spontaneously was discussed during the importance of development technics to reuse the water, a important sustainability topic.
We talked about how the Artificial Intelligence would be used to monitor the body condition during an extraterrestrial travel: controlling some health levels of the space travellers.
We were sleep after midnight, reflecting about many interesting, not only space, but human topics, such as consciousness: fundamental subject for the sustainability of space missions.
I consider that in Space Mission and in Mars missions would be necessary some time to be flexible to think creatively. So, creativity and innovation is the key to We arrive in new worlds. Our experience in MDRS is a good place to think out of the box, about this advanced topics.
Everybody wake up later. No stress! J Today We not had Extravehicular Activities (EVAs). So, some of us explored some possibilities od the MDRS facilities and develop new tasks related to their research.
8h42m – 9h17m – I walked and ran during 35 minutes in the internal corridor of the MDRS, a distance equivalent to 3,5km or more than 2 miles. I consider do more of this. Before the mission was walking and running between 1 and 2,5 hours. I consider that physical activities it is also important to help us to think more creatively. I am motivated to encourage my crewmates also do some exercise during this time.
One of the topics of my research is to evaluate how to deal with distance work activities. During the first days I had a lot of duties. So, today, during the morning, I presented a time to contact my partners in Brazil and also paid some bills using the internet banking services. In Mars this also will an important issue, because a person in mars continues to have some obligations in the planet Earth. Later our router was turned off. In Mars will be also times without communication with Earth. So, this is also part of our simulation.
Also during the morning, I applied my survey related with sustainability with our crew. In the end of the mission I will apply a new survey and compare the result the perception of some sustainability issues perception that can be changed during the 2 weeks mission.
In the afternoon, Atila started to do his agriculture experiment, comparing the use of regolith and the regular soil. The soils will be used to grow seeds and than He will compare the results generated by the two different soils.
I had prepare some meals for the crew. I consider that a very important issue of the life in a habitat Mars related is the wellbeing. The first missions to Mars will be so unique and expensive that all the team needs to arrive will health in Mars. In the planet, the daily life is also important. I am, as a commander, consider be my responsibility to promote a better experience for all the crew.
Each one presents their own psychological, social and physical space, and this needs to be respected and identify the best contribution of each. We are a great team!
13h40m – I started to prepare the lunch.
14h30m – Our lunch was: spaghetti with tuna and cheese, tomato soup, peas and broccoli. After the lunch We discussed some topics. Part of the crew still worried about the internet: if We will consume all the data available, etc. Part of the day the modem will be kept turn off.
16h35m – I walked more 25 minutes, completing 1 hour today. According to Apple Watch I walked 7,8 km or almost 5 miles achieving health goals for today! J I will try to keep those exercises. It perceivable how cold is outside today.
17h30m – I’m preparing a book about many aspects related with Mars, MDRS and some experiences that I’m working in Brazil: the research station Habitat Marte (www.HabitatMarte.Blogspot.com),
Informamos que foram selecionados os seguintes candidatos para participar da primeira missão na Estação de Pesquisa Habitat Marte, a ocorrer no período de 48 …
MarsLab – Laboratory of Technologies Applied to Mars and Semiarid (www.MarsLaboratory.Blogspot.com), Mars Research Group – Mars and Arid Regions Settlements Sustainability and the Museu Espaço Marte (www.espacoMarte.Blogspot.com). I’m also committed to create the Mars Society Brazil chapter.
|Museu Espaço Marte
O Museu Espaço Marte é um ambiente educativo que tem como propósito apresentar a cultura espacial com foco especial em Marte, com associação ao desenvolvimento …
18h15m –Tonight I review all the MDRS Forms Instruction.
19h – CapComm.
MDRS Daily Summary Report for sol
Stinky like Pepe LePew but na, na, na, stayin’ alive!
Mission Status: We’re OK. There are no incidents to report.
Sol Activity Summary:
First SIM day went pretty well. Our breakfast was simple enough and then, we prepared for the EVA session. There is a lot going on in those sessions, though. We had some disgareements at first, but everything went well at 2 p.m. (EVA time!) There was some trouble with the radios, though, but we will try to fix that tonight. Weather was fine, too, although the wind decided to blow with some power, just in our EVA session! But we’re pretty awesome, so our capacity to resist is above average.
The crew scientists are being pretty grumpy at this hour of the night. Some are worried about their plant pals! That doesn’t usually happenon Earth! Is this a consequence of Martian issolation? Who knows?
Look Ahead Plan: Tomorrow, we’ll continue with our experiments. It will be a hab day, tomorrow. Orobably we will repare some things and rearrange the lab. It needs a serious re-decoration. The Green Hab will need some repairs too.
Anomalies in work: We just had troubles with the radios, really. But we will figure something out later tonight.
Weather: … Windy. Somewhat cloudy and just chilly.
Crew Physical Status: Everyone is in good shape. Thanks to our everyday meals! There is some tiredness, though, thanks to the EVA.
EVA: We just explored around the hab, less than a mile.
Reports to be file: Already sent. Please tell us if you need everything in one mail.
Support request: Fire Alarms. Better explained at the HSO officer. Light for the Lab and Green Hab.
Camila Castillo, Atila Meszaros
Crew Scientist, Green Hab Officer
Mars is doing pretty fine right now. As the Crew Scientist, I’ll report the principal events of today’s experiments:
Yesterday, we started the germination tests for the raddish seeds. If everything goes ok, then the seeds will germinate in the next 3 or 4 days. This will be necessary for determinating if these seeds can grow.
We recollected soil samples in today’s EVA. It should be enough for one phase of the geology experiment and one phase of the two biology experiments.
The lab is ok, too. We started to relocate some stuff in there, also labeling it. We want to be able to find all the stuff whenever we need it.
We still can’t report any results, though, but hopefully we will have something for the first days of the next week.
As the Crew Journalist, it seems that I will not work that much today. However, I’ll be writing some stuff about todays events in both english and spanish, so please tell me if you would like some of this writings.
As I told the Crew, I love to write, so I don’t have any problems telling jokes via e-mail!
Hope everything is great on Earth!
Crew Scientist and Journalist
A great challenge for me is to prepare the meals in the first days of our simulation. In the first days, I considered had did delicious meals. I never had prepare a meal as I had done here. So, for me, it is a great learning that I will bring back to Brazil. Thank you MDRS and Mars Society for makes emerges in me some new competencies applied to my life. J
7h45m – I spent 1 hour walking in the MDRS corridors, in a distance of 6,4 km or 4 miles.
8h55m – In Hanksville is doing 1oC. or 34oF.
9h – Preparation of the breakfast.
10h30m – We still in breakfast. One of the main topics of our debates is related to internet. Just in the first hours, our internet just consumed 290 MB from 550 MB available. So, We are very worried about that. K
10h45m – My soundtrack today is Megadeth album Symphony of Destruction during the time when I’m writing the book related to Mars and Sustainability.
12h – 13h – Shannon Rupert (MDRS Director) presented to us a very detailed instruction about the use of the All-Terrain Vehicles (ATV). Shannon is always very friendly with us.
13h – Start to prepare the lunch.
13h20m – Stop to help in the EVA preparation.
13h30m – Atila, Brandon and Carmen develop an EVA activity using the ATV. The returned 15h30m.
13h40m – Return to prepare the lunch.
14h40 – Just preparing the lunch.
15h50m-16h50m – Lunch. Our meal today was brown beans, rice with broccoli and beef dehydrated.
17h20m – I just walked more 30 minutes of 4,8 km or 3 miles, completing 11,2 km of walking, or 7 miles.
We all considered that the day was so short today and the time pass so fast.
Prof. Julio Rezende, PhD
MDRS Daily Summary Report for sol 5
Summary Title: Time is running fast today.
Mission Status: The day was very busy today. All the crew members had something to do.
Sol Activity Summary: Prepare the meal today was so painful. I never cooked beans. It looks like old. I had some problem with it. Our crew had a good time today. All doing different and complementary activities. We are all well motivated and keeping a good relationship. Thank You for all presenting the wishes relate with us.
Look Ahead Plan: Tomorrow, We are waiting the Brazilian reporters of TV Globo at 9 AM. We are planning an EVA during their time here. We are just prepared to receive them.
Anomalies in work: Nope.
Weather: 6 C, 43 F
Crew Physical Status: Fine.
EVA: today happened EVA with the ATV.
Reports to be file: We sent in CapComm.
Support Requested: None.
Mars today has been really great, with bright sunshine and now a cozy hab to sleep in!
Things in the lab have been really great! The germination tests of the raddish seeds are going great. 3/5 in one plate and 3/5 in the second plate have presented a positive response, so the seeds are viable. Raddish seeds are also really fast at germinating, but this new brand we bought is particullarly fast. This will be great for the lab tests.
The new autoclave is now at the lab, and we have plenty of destillated water, so that will help a lot with the different biological experiments. The new microscope is also set up and ready to go! It also incorporates a new cam for our sights in it, so I would like to ask where should I install the drive to use the cam. I fear that this drive is just a one-time software, so if I install it in my computer, it would be the only time it could be installed. That happens quite a lot with other copyrighted softwares. I tried to install it in the hab laptop, but we couldn’t find the cable to connect it.
At today’s EVA, our crew as able to recollect more soil for their experiments. A crewmate is trying to tipify the soil, while other project involves the germination of different seeds in two types of soil.
There’s also this idea some of us had of using the soapy water for some seeds. I’m running a pilot test with raddish seeds. It would be interesting if we find a way of making that dirty soapy water from the dishes useful!
I’ll attach some photos of the experiments in the lab today. The first two ones are from my raddish experiments. The last ones are from the pilot experiment with the dirty water from the dishes.
We also tried to organize the lab better. There are now labels in each drawer and we are trying to set places for the equipment. There is a table for running microbiological experiments and another for geological ones. We are hoping to make the lab a great and ordered space for future crews!
Non-nominal systems: Working
Solar: For now, not given power to the Hab,
Diesel : 64%
Propane – 62psi
Gasoline (5 Gallon containers for ATV) – 0 gallons
Water (trailer) – empty
Water (static) –
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 519 gallons
ATV’s Used: Yes
Oil Checked: motos y carros: Yes
Atv # Fuel Used Gals: 2 gallons
Tires Status: Perfect
# Hours Used Day: 4hours
Phobos used: No
Beginning charge: –
Ending charge: –
Currently charging: –
Deimos used: Yes
Beginning charge: 4pm
Ending charge: Continuing charging
Currently charging: Yes
HabCar used?: No
Notes and Comments:
Last night was a little cold, we are having problem with the heater., because when we turn it on, it heats, but few hours later the air is cold again. So, we have to turn it off and then turn it on again to keep us warm. Do you know a solution for this problem?
At the same time, we did an EVA session today. The blue moto is not working. In that way, we did the EVA session with the deimo and the red moto. We do not had any problems with these vehicles. In contrast, we had problems with the radios. All radios are working, however, approximately 2.5 km far from the Hab, we could not had any communication with the Hab. We do not know why that is. So, please send us the ratio of distance in which we could keep the communication.
Tomorrow we are going to use the solar.
We started EVA at 1:35 PM, instead of the planned 1:00 PM. The delay was due to an extended period for our ATV training to take place. The EVA crew included myself, Carmen, and Atila. Drove north and spotted vegetation growth along road and marked the location. We tried marking the location with a GPS, but we found it was impossible to read off the screen with the helmets on, since the screen was too scratched up. Instead, I determined the coordinates by closely examining the planning map after the EVA. We Took several pictures of the area along with the surrounding terrain
We continued north and passed Tank Wash due to navigational Error. We ended up taking the road all the way to the Quarry Site. We retraced our steps southward until we found the reservoir dam, which was the most distinguishable feature I could use to place our location.
-Stopped at coordinates:
We then slowly drove northward, until we found the correct wash
-Stopped at coordinates:
We made our way to a talus slope that contained broken down rock from the cliff above us. Upon digging through the cemented layers, we found three distinct layers within only 20 cm below the surface.
– Coordinates of Sample Location #3:
There was a white-colored layer in the first 7 cm, then a green layer below it that was 3-4 cm thick. Below that was a dark brown / black soil lightly peppered with white grains. We collected samples from each of these layers and photographed the area and our activities. We loaded up our samples and returned to the hab. Before we entered the airlock, Carmen checked the Diesel and Propane tank levels. We began recompression in the airlock by 3:45 PM.
For the next EVA, we need to improve our radio Comms. I’ll instruct the crew members on words and phrases to use on the radio, such as “affirmative”, “negative”, “over”, “loud and clear”, etc. Also going over the phonetic alphabet may come in handy. In order to reduce the chance of another navigational error, I’ll have to take very special notice of the surface features surrounding my intended destination before we depart. I’d also like to request obtaining another GPS device that we can read on the field so we can record our sample locations in the future.